Express 7920 in radians and (7π/12)c in degrees.
Answer by stanbon(75887)

Express 7920 in radians and (7π/12)c in degrees.
--------
Since 180 degrees = pi radians, (180/pi) converts radians to degrees
and (pi/180) converts degrees to radians.
------
7920(pi/180) = 44pi radians
------
(7pi/12)(180/pi) = 7*15 = 105 degrees
======================================
Cheers,
Stan H.
========================